Find the the equation of a circle which passes through the intersection points of the the lines x - y - 3 = 0 and 3x - 10y + 5 = 0 and centered at ( 2, - 2 ).

Grade:11

1 Answers

Piyush Kumar Behera
7 years ago
 IS (5,2) .intersection points of the the lines x - y - 3 = 0 and 3x - 10y + 5 = 0 
RADIUS = DISTANC BETWEEN THE POINTS (2,- 2) AND  (5,2) 
           =\sqrt{9+16} =5
EQUATION OF THE CIRCLE=(X-2)2+(Y+2)2=25